# FG-CLIP 2: A Bilingual Fine-grained Vision-Language Alignment Model

### v2

@@ Abstract
Fine-grained vision-language understanding requires precise alignment between visual content and linguistic descriptions, a capability that remains limited in current models, particularly in non-English settings. While models like CLIP perform well on global alignment, they often struggle to capture fine-grained details in object attributes, spatial relations, and linguistic expressions, with limited support for bilingual comprehension. To address these challenges, we introduce FG-CLIP 2, a bilingual vision-language model designed to advance fine-grained alignment for both English and Chinese. Our approach leverages rich fine-grained supervision, including region-text matching and long-caption modeling, alongside multiple discriminative objectives. We further introduce the Textual Intra-modal Contrastive (TIC) loss to better distinguish semantically similar captions. Trained on a carefully curated mixture of large-scale English and Chinese data, FG-CLIP 2 achieves powerful bilingual performance. To enable rigorous evaluation, we present a new benchmark for Chinese multimodal understanding, featuring long-caption retrieval and bounding box classification. Extensive experiments on 29 datasets across 8 tasks show that FG-CLIP 2 outperforms existing methods, achieving state-of-the-art results in both languages. We release the model, code, and benchmark to facilitate future research on bilingual fine-grained alignment.
**footnotetext: Equal contribution. E-mail: xiechunyu@360.cn, wangbin10@360.cnfootnotetext: $\dagger$} Corresponding Author. E-mail: lengdawei@360.cn
@@ 1 Introduction
Several recent works have sought to address these limitations. Approaches such as FineCLIP (Jing et al., 2024) and LongCLIP (Zhang et al., 2024a) improve fine-grained understanding by incorporating region-level signals or supporting longer textual inputs. FG-CLIP (Xie et al., 2025) significantly advances the state of fine-grained modeling through large-scale data curation and attribute-based hard negative sampling. On the language side, Chinese-CLIP (Yang et al., 2022) and R2D2 (Xie et al., 2023) have laid the groundwork for Chinese vision-language alignment, yet they primarily focus on short-caption retrieval and lack support for fine-grained or region-level tasks. The development of fine-grained capabilities has thus remained largely confined to English, whereas Chinese models operate at a coarser semantic level. No existing framework unifies these directions, and there is a notable absence of comprehensive benchmarks for evaluating fine-grained understanding in Chinese, which hinders systematic progress in bilingual multimodal research.
@@
We further contribute a new benchmark suite to advance evaluation in Chinese multimodal understanding, featuring challenging tasks such as long caption image-text retrieval and bounding box classification in Chinese that go beyond conventional short-text retrieval and assess fine-grained comprehension more rigorously. Extensive experiments show that FG-CLIP 2 outperforms existing models on 29 datasets across 8 vision-language tasks in both Chinese and English, demonstrating powerful bilingual fine-grained vision-language alignment capability. To support future research and real-world deployment, our model, code and benchmark will be made publicly available.
@@ 2 Related Work
Vision-language alignment models trained on large-scale data, such as CLIP (Radford et al., 2021), EVA-CLIP (Sun et al., 2023), SigLIP (Zhai et al., 2023), MetaCLIP (Xu et al., 2024) and DFN (Fang et al., 2024) have demonstrated strong zero-shot capabilities but primarily focus on global semantic alignment and are often trained on English-only corpora. While these models serve as backbones for downstream tasks including multimodal reasoning (Bai et al., 2025), open-vocabulary detection (Fu et al., 2025), and segmentation (Cho et al., 2024; Wang et al., 2025b), their lack of fine-grained and multilingual understanding limits broader applicability.
Recent efforts aim to improve localization capability and dense feature alignment through region-level supervision. Methods like AlphaCLIP (Sun et al., 2024), FineCLIP (Jing et al., 2024), and FG-CLIP (Xie et al., 2025) leverage bounding boxes or masked regions to enhance local correspondence, while CLOC (Chen et al., 2024a), TIPS (Maninis et al., 2025), and SigLIP 2 (Tschannen et al., 2025) introduce architectural or training enhancements for richer feature generation. On the multilingual front, Chinese-CLIP (Yang et al., 2022) and R2D2 (Xie et al., 2023) target Chinese understanding, and MetaCLIP 2 (Chuang et al., 2025) scales multilingual data collection for broader language coverage. However, these works often treat fine-grained and multilingual understanding separately, and none explicitly optimize both in a unified framework. This gap inspires our work.
Figure 1: Overview of the FG-CLIP 2.
@@ 3.1 Architecture
We build upon the SigLIP 2 (Tschannen et al., 2025) dual-encoder framework, introducing key adaptations for fine-grained understanding and bilingual alignment. For the text encoder, we extend the maximum input length from 64 to 196 tokens to accommodate longer descriptions. We employ the multilingual Gemma tokenizer (Team et al., 2024) with a 256K vocabulary. On the vision side, we adopt a data-adaptive resolution strategy: the target resolution is selected from {128, 256, 576, 784, 1024} based on the maximum image size per mini-batch, avoiding the stochastic sampling of SigLIP 2 and ensuring consistent training and inference behavior with minimal upscaling or downscaling. Vision and text features are aggregated using a masked attention pooling (MAP) head (Zhai et al., 2022). We implement three model variants (Base, Large, and So400M (Alabdulmohsin et al., 2023)) to evaluate performance across different scales.
@@ 3.2 Training Objectives
Our training proceeds in two stages. In Stage I, we optimize only the global alignment objective. In Stage II, we jointly optimize five complementary objectives:
$\mathcal{L}=\lambda_{1}\mathcal{L}_{\text{Global}}+\lambda_{2}\mathcal{L}_{\text{FGV}}+\lambda_{3}\mathcal{L}_{\text{FGT}}+\lambda_{4}\mathcal{L}_{\text{CMR}}+\lambda_{5}\mathcal{L}_{\text{TIC}},$ (1)
with fixed weights $\lambda_{1}=1.0$, $\lambda_{2}=0.1$, $\lambda_{3}=0.5$, $\lambda_{4}=0.4$, $\lambda_{5}=0.1$, chosen to ensure stable and effective multi-objective optimization. The following paragraphs detail each component.
Global Alignment Learning.
We adopt the sigmoid loss from SigLIP (Zhai et al., 2023), which treats image-text matching as a binary classification. For each image-text pair, similarity scores are computed across all pairs in the batch, and logistic regression is applied to distinguish positive from negative pairs. To enrich textual supervision, we include both original short captions and long captions generated by Large Multimodal Models (LMMs). This dual-caption strategy provides complementary signals: concise labels for global semantics and detailed descriptions for richer context.
Fine-Grained Visual Learning.
To enable sub-image alignment, we generate dense image embeddings by introducing an additional self-attention module, circumventing the information bottleneck caused by CLS-only pooling in SigLIP 2. For each annotated region, we extract patch-level embeddings and apply RoIAlign (He et al., 2017) to obtain region-specific features. Corresponding text embeddings are derived from phrase-level descriptions aligned to each bounding box. The regional contrastive loss encourages alignment between matched region-text pairs, promoting fine-grained cross-modal understanding.
Fine-Grained Textual Learning.
Following FG-CLIP (Xie et al., 2025), we leverage hard negatives from the FineHARD dataset. Each positive region-text pair is paired with 10 semantically similar negatives, constructed by perturbing key attributes (e.g., color, count, action) while preserving syntactic structure. The model is trained with a binary classification loss over the 1 positive and 10 negatives, encouraging the model to assign higher scores to the matched pair. This objective enhances the model’s ability to discriminate subtle textual differences.
Cross-modal Rank Loss with Global Threshold Synchronization.
We adopt the Cross-modal Rank (CMR) loss (Zhang et al., 2024b) to enforce a margin between positive and hard negative pairs, thereby strengthening the model’s discrimination of semantic boundaries. For a positive pair $(I,T)$ and hard negative $T_{k}$, the loss is defined as:
@@
Textual Intra-modal Contrastive Loss.
While cross-modal alignment ensures image-text correspondence, the text encoder itself often lacks sufficient discriminative pressure to separate semantically similar but distinct region descriptions, which is critical for fine-grained visual grounding. To address this, we propose the Textual Intra-modal Contrastive (TIC) loss, which operates purely within the text modality to sharpen the text encoder’s representation space. Given a batch of region texts, we compute pairwise similarities and filter out pairs with $\text{sim}>0.95$ to avoid over-penalization. The top-10 most similar texts per sample are selected as hard negatives. The TIC loss is then defined as:
$\mathcal{L}_{\text{TIC}}=-\sum_{i=1}^{N}\log\frac{1}{\sum_{T_{m}\in\mathcal{T}_{i}}{\mathrm{exp}}(S(T_{i},T_{m}))},$ (4)
@@ 3.3 Training Data
In the first stage, we train on image-text pairs from diverse sources, with a particular emphasis on enhancing semantic depth and linguistic coherence. For English, we adopt an enhanced version of the LAION-2B dataset (Xie et al., 2025), where we augment the original short captions with detailed long captions generated by LMMs. The original captions are often fragmented and contain keyword stacking or irrelevant noise, making them insufficient for training models to understand rich, compositional language. We retain original captions to preserve diversity of natural language expressions, while simultaneously training on their enhanced long-caption counterparts. This dual-caption strategy enables the model to learn from both concise, real-world descriptions and semantically dense, contextually coherent narratives. For Chinese, we combine three datasets: Wukong (100 million pairs) (Gu et al., 2022), Zero (250 million pairs) (Xie et al., 2023), and a large-scale in-house dataset (500 million pairs).
In the second stage, we extend training with fine-grained region-text pairs to further improve spatial grounding. For English, we use the FineHARD dataset (Xie et al., 2025), which includes 12 million images, 40 million bounding boxes with fine-grained region descriptions, and 10 million hard negative samples. For Chinese, we use an in-house dataset containing 12 million images. An overview of our training datasets is provided in Table A in Appendix.
@@ 3.4 Bilingual Evaluation Protocols
Existing multimodal benchmarks in English are diverse and well-established, covering a broad spectrum of vision-language tasks such as fine-grained understanding (e.g., FG-OVD (Bianchi et al., 2024)), open-vocabulary object detection (e.g., LVIS (Gupta et al., 2019), COCO (Lin et al., 2014)), and image-text retrieval (e.g., Flickr30K (Young et al., 2014), DCIUrbanek et al. (2024)). These resources enable comprehensive evaluation of model capabilities across different granularities and semantic complexities. In contrast, Chinese multimodal datasets remain limited in scope and diversity, with most focusing on short-caption retrieval tasks such as COCO-CN (Li et al., 2019) and Flickr30k-CNA (Xie et al., 2023). Such benchmarks are insufficient for evaluating fine-grained cross-modal alignment, particularly at the region level or with long, descriptive textual inputs.
Table 1: Performance comparison on fine-grained understanding and bounding box classification tasks using Top-1 accuracy.
Method Backbone Fine-Grained Understanding BBox Classification
Hard Medium Easy Trivial COCO80 LVIS1203 BoxClass-CN566
CLIP ViT-B/16 12.0 23.1 22.2 58.5 44.2 20.9 -
EVA-CLIP ViT-B/16 14.0 30.1 29.4 58.3 30.6 14.4 -
Long-CLIP ViT-B/16 9.2 18.4 16.2 51.8 36.7 18.2 -
FineCLIP ViT-B/16 26.8 49.8 50.4 71.9 48.4 23.3 -
SigLIP 2 ViT-B/16 24.9 46.5 48.7 85.0 53.4 20.6 57.9
FG-CLIP ViT-B/16 46.1 66.6 68.7 83.4 52.3 28.6 -
FG-CLIP 2 ViT-B/16 52.3 76.3 80.3 92.0 74.9 47.3 60.7
CLIP ViT-L/14 15.4 25.3 25.7 38.8 33.8 9.3 -
EVA-CLIP ViT-L/14 18.3 38.4 35.2 62.7 32.1 18.3 -
Long-CLIP ViT-L/14 9.6 19.7 16.0 39.8 35.6 10.4 -
FineCLIP ViT-L/14 22.8 46.0 46.0 73.6 54.5 22.5 -
SigLIP 2 ViT-L/16 24.1 47.1 47.4 84.1 54.7 25.9 56.6
FG-CLIP ViT-L/14 48.4 69.5 71.2 89.7 63.2 38.3 -
FG-CLIP 2 ViT-L/16 55.5 77.5 83.1 92.5 74.0 41.9 68.6
Meta CLIP 2 ViT-H/14 16.5 36.6 34.7 79.6 52.0 24.4 55.2
SigLIP 2 ViT-So/16 26.0 48.7 49.9 87.4 62.0 31.4 63.6
FG-CLIP 2 ViT-So/16 54.0 77.4 79.8 93.5 77.4 43.9 66.5
@@
BoxClass-CN is a region classification dataset that evaluates the alignment between image regions and their corresponding Chinese textual descriptions. It complements existing Chinese benchmarks by providing region-level supervision and serves as an evaluation suite for assessing models’ fine-grained understanding of visual content. We construct this dataset through a scalable automated pipeline based on the LAION-2B corpus (Schuhmann et al., 2022). We first sample 200,000 images and generate detailed captions using a LMM (Hong et al., 2024). These captions are parsed to extract referring expressions, which are then localized using a pretrained object detector (Cheng et al., 2024) to produce bounding box proposals. Non-maximum suppression removes overlapping boxes, and only those with region-text similarity above 0.15 (computed by FG-CLIP (Xie et al., 2025)) are retained. Detected categories undergo semantic clustering and merging, resulting in 566 semantically refined categories. These categories are translated into Chinese, and the final dataset consists of 24,221 images and 66,258 high-quality region-text pairs.
Together, these datasets provide a more rigorous and comprehensive assessment for bilingual vision-language alignment models, supporting deeper evaluation of fine-grained understanding capability.

### v3

@@ Abstract
Fine-grained vision-language understanding requires precise alignment between visual content and linguistic descriptions, a capability that remains limited in current models, particularly in non-English settings. While models like CLIP perform well on global alignment, they often struggle to capture fine-grained details in object attributes, spatial relations, and linguistic expressions, with limited support for bilingual comprehension. To address these challenges, we introduce FG-CLIP 2, a bilingual vision-language model designed to advance fine-grained alignment for both English and Chinese. Our approach leverages rich fine-grained supervision, including region-text matching and long-caption modeling, alongside multiple discriminative objectives. We further introduce the Textual Intra-modal Contrastive (TIC) loss to better distinguish semantically similar captions. Trained on a carefully curated mixture of large-scale English and Chinese data, including a newly released 12M Chinese region-text dataset, FG-CLIP 2 achieves powerful bilingual performance. To enable rigorous evaluation, we present a new benchmark for Chinese multimodal understanding, featuring long-caption retrieval and bounding box classification. Extensive experiments on 29 datasets across 8 tasks show that FG-CLIP 2 outperforms existing methods, achieving state-of-the-art results in both languages. We release the model, code, and benchmark to facilitate future research on bilingual fine-grained vision-language alignment.
Machine Learning, ICML
Homepage: https://360cvgroup.github.io/FG-CLIP
Code: https://github.com/360CVGroup/FG-CLIP
Model&Dataset: https://huggingface.co/collections/qihoo360/fg-clip-2
@@ 1 Introduction
Several recent works have sought to address these limitations. Approaches such as FineCLIP (Jing et al., 2024) and LongCLIP (Zhang et al., 2024a) improve fine-grained understanding by incorporating region-level signals or supporting longer textual inputs. FG-CLIP (Xie et al., 2025) further advances fine-grained discrimination through large-scale data curation and attribute-aware hard negative sampling. However, achieving robust discrimination between highly similar textual descriptions or subtle visual variations remains challenging, suggesting opportunities for more advanced training objectives. Moreover, prior fine-grained approaches are predominantly developed and evaluated in English, leaving Chinese–English bilingual scenarios relatively underexplored. Meanwhile, models designed for Chinese, such as Chinese-CLIP (Yang et al., 2022) and R2D2 (Xie et al., 2023), primarily focus on coarse-grained alignment and lack fine-grained modeling capability. Therefore, extending fine-grained alignment to bilingual vision–language settings remains an important yet underexplored problem.
@@
We further contribute a large-scale Chinese region-text dataset, FineRegion-CN, containing 12 million images with fine-grained region descriptions to fill the gap of Chinese region-level training data. Additionally, we introduce a new benchmark suite to advance evaluation in Chinese multimodal understanding, featuring challenging tasks such as long caption image-text retrieval and bounding box classification in Chinese that go beyond conventional short-text retrieval and assess fine-grained comprehension more rigorously. Extensive experiments show that FG-CLIP 2 outperforms existing models on 29 datasets across 8 vision-language tasks in both Chinese and English, demonstrating powerful bilingual fine-grained vision-language alignment capability. To support future research and real-world deployment, our model, training data, code, and benchmark are made publicly available.
@@ 2 Related Work
Vision-language alignment models trained on large-scale data, such as CLIP (Radford et al., 2021), EVA-CLIP (Sun et al., 2023), SigLIP (Zhai et al., 2023), MetaCLIP (Xu et al., 2024) and DFN (Fang et al., 2024) have demonstrated strong zero-shot capabilities but primarily focus on global semantic alignment and are often trained on English-only corpora. While these models serve as backbones for downstream tasks including multimodal reasoning (Bai et al., 2025), open-vocabulary detection (Fu et al., 2025), and segmentation (Cho et al., 2024; Wang et al., 2025b), they lack fine-grained and multilingual understanding, limiting their broader applicability.
Recent efforts aim to improve localization capability and dense feature alignment through region-level supervision. Methods like AlphaCLIP (Sun et al., 2024), UMG-CLIP (Shi et al., 2024), FineCLIP (Jing et al., 2024), and FG-CLIP (Xie et al., 2025) leverage bounding boxes or masked regions to enhance local correspondence, while CLOC (Chen et al., 2024a), TIPS (Maninis et al., 2025), and SigLIP 2 (Tschannen et al., 2025) introduce architectural or training enhancements for richer feature generation. On the multilingual front, Chinese-CLIP (Yang et al., 2022) and R2D2 (Xie et al., 2023) target Chinese understanding, and MetaCLIP 2 (Chuang et al., 2025) scales multilingual data collection for broader language coverage. However, these works often treat fine-grained and multilingual understanding separately, and none explicitly optimize both in a unified framework. This gap inspires our work.
Figure 1: Overview of FG-CLIP 2. The framework consists of a two-stage training pipeline. Stage I performs bilingual global alignment using data-adaptive image resolution, matching image global tokens with long and short caption embeddings via $\mathcal{L}_{\text{Global}}$. Stage II introduces fine-grained region-caption supervision alongside the global alignment. It employs ROI features and explicit hard-negative mining to compute auxiliary objectives: Fine-Grained Visual Learning ($\mathcal{L}_{\text{FGV}}$), Fine-Grained Textual Learning ($\mathcal{L}_{\text{FGT}}$), Textual Intra-modal Contrastive loss ($\mathcal{L}_{\text{TIC}}$), and Cross-modal Rank Loss ($\mathcal{L}_{\text{CMR}}$), thereby enhancing discrimination for semantic details.
@@ 3.1 Architecture
We build upon the SigLIP 2 (Tschannen et al., 2025) dual-encoder framework, introducing key adaptations for fine-grained understanding and bilingual alignment. For the text encoder, we extend the maximum input length from 64 to 196 tokens to accommodate longer descriptions. On the vision side, we adopt a data-adaptive resolution strategy: the target resolution is selected from {128, 256, 576, 784, 1024} based on the maximum image size per mini-batch, avoiding the stochastic sampling of SigLIP 2 and ensuring consistent training and inference behavior with minimal upscaling or downscaling. Global vision features are aggregated via a masked attention pooling head (Zhai et al., 2022), while region embeddings are generated through an additional self-attention module. This design enables fine-grained, sub-image alignment and circumvents the information bottleneck caused by CLS-only pooling.
@@ 3.2 Training Objectives
Our training proceeds in two stages. Stage I focuses on Global Alignment Learning. In Stage II, this objective is jointly optimized with four additional learning objectives, detailed in the subsequent paragraphs.
Global Alignment Learning. We adopt the sigmoid loss from SigLIP (Zhai et al., 2023), which treats image-text matching as a binary classification. For each image-text pair, similarity scores are computed across all pairs in the batch, and logistic regression is applied to distinguish positive from negative pairs. To enrich textual supervision, we include both original short captions and long captions generated by Large Multimodal Models (LMMs). This dual-caption strategy provides complementary signals: concise labels for global semantics and detailed descriptions for richer context. We denote this global alignment loss as $\mathcal{L}_{\text{Global}}$.
Fine-Grained Visual Learning. For each annotated region, we extract patch-level embeddings and apply RoIAlign (He et al., 2017) to obtain region-specific features. Corresponding text embeddings are derived from descriptions aligned to each bounding box. The regional contrastive loss encourages alignment between matched region-text pairs, promoting fine-grained cross-modal understanding. This loss term is denoted as $\mathcal{L}_{\text{FGV}}$.
Fine-Grained Textual Learning. Following FG-CLIP (Xie et al., 2025), we leverage hard negatives from the FineHARD dataset. Each positive region-text pair is paired with 10 semantically similar negatives, constructed by perturbing key attributes (e.g., color, count, action) while preserving syntactic structure. The model is trained with a binary classification loss over the 1 positive and 10 negatives, encouraging the model to assign higher scores to the matched pair. This objective enhances the model’s ability to discriminate subtle textual differences. We formulate this discriminative objective as $\mathcal{L}_{\text{FGT}}$.
Textual Intra-modal Contrastive Loss. While cross-modal alignment ensures image-text correspondence, the text encoder itself often lacks sufficient discriminative pressure to separate semantically similar but distinct region descriptions, which is critical for fine-grained visual grounding. To address this, we propose the Textual Intra-modal Contrastive (TIC) loss, which operates purely within the text modality to sharpen the text encoder’s representation space. Given a batch of region texts, we compute pairwise similarities and filter out pairs with $\text{sim}>0.95$ to avoid over-penalization. The top-10 most similar texts per sample are selected as hard negatives. The TIC loss is then defined as:
$\mathcal{L}_{\text{TIC}}=-\sum_{i=1}^{N}\log\frac{1}{\sum_{T_{m}\in\mathcal{T}_{i}}\exp(S(T_{i},T_{m}))},$ (1)
Cross-modal Rank Loss with Global Threshold Synchronization. We adopt the Cross-modal Rank (CMR) loss (Zhang et al., 2024b) to enforce a margin between positive and hard negative pairs, thereby strengthening the model’s discrimination of semantic boundaries. For a positive pair $(I,T)$ and hard negative $T_{k}$, the loss is defined as:
@@
In Stage II, the training objective combines these terms as:
$\mathcal{L}=\mathcal{L}_{\text{Global}}+\lambda_{1}\mathcal{L}_{\text{FGV}}+\lambda_{2}\mathcal{L}_{\text{FGT}}+\lambda_{3}\mathcal{L}_{\text{TIC}}+\lambda_{4}\mathcal{L}_{\text{CMR}},$ (4)
with fixed weights $\lambda_{1}=0.1$, $\lambda_{2}=0.5$, $\lambda_{3}=0.1$, $\lambda_{4}=0.4$, chosen to ensure stable and effective multi-objective optimization. The selection strategy for loss weights is detailed in Appendix F.
@@ 3.3 Training Data
In the first stage, we train on image-text pairs from diverse sources, with a particular emphasis on enhancing semantic depth and linguistic coherence. For English, we adopt an enhanced version of the LAION-2B dataset (Xie et al., 2025), which augments the original short captions with detailed long captions generated by LMMs. The original captions are often fragmented and contain keyword stacking or irrelevant noise, making them insufficient for training models to understand rich, compositional language. We retain original captions to preserve diversity of natural language expressions, while simultaneously training on their long-caption counterparts. This dual-caption strategy enables the model to learn from both concise, real-world descriptions and semantically dense, contextually coherent narratives. For Chinese, we combine three datasets: Wukong (100M pairs) (Gu et al., 2022), Zero (250M pairs) (Xie et al., 2023), and a large-scale in-house dataset (500M pairs).
In the second stage, we extend training with fine-grained region-text pairs to further improve spatial grounding. For English, we use the FineHARD dataset (Xie et al., 2025), which includes 12 million images, 40 million bounding boxes with fine-grained region descriptions, and 10 million hard negative samples. For Chinese, we construct a complementary dataset, termed FineRegion-CN, following the region-text annotation pipeline of FineHARD. It contains 12 million images with bounding boxes and fine-grained region descriptions in Chinese. FineRegion-CN are publicly released to support research on Chinese fine-grained multimodal understanding. An overview of our training datasets is provided in Appendix A.
@@ 3.4 Bilingual Evaluation Protocols
Existing multimodal benchmarks in English are diverse and well-established, covering a broad spectrum of vision-language tasks such as fine-grained object-level understanding (e.g., FG-OVD (Bianchi et al., 2024)), open-vocabulary object detection (e.g., LVIS (Gupta et al., 2019), COCO (Lin et al., 2014)), and image-text retrieval (e.g., Flickr30K (Young et al., 2014), DCI(Urbanek et al., 2024)). These resources enable comprehensive evaluation of model capabilities across different granularities and semantic complexities. In contrast, Chinese multimodal datasets remain limited in scope and diversity, with most focusing on short-caption retrieval tasks such as COCO-CN (Li et al., 2019) and Flickr30k-CNA (Xie et al., 2023). Such benchmarks are insufficient for evaluating fine-grained cross-modal alignment, particularly at the region level or with long, descriptive textual inputs.
@@
BoxClass-CN is a region classification dataset that evaluates the alignment between image regions and their corresponding Chinese textual descriptions. It complements existing Chinese benchmarks by providing region-level supervision and serves as an evaluation suite for assessing models’ fine-grained understanding of visual content. We construct this dataset through a scalable automated pipeline based on the LAION-2B corpus (Schuhmann et al., 2022). We first sample 200,000 images and generate detailed captions using a LMM (Hong et al., 2024). These captions are parsed to extract referring expressions, which are then localized using a pretrained object detector (Cheng et al., 2024) to produce bounding box proposals. Non-maximum suppression removes overlapping boxes, and only those with region-text similarity above 0.15 (computed by FG-CLIP (Xie et al., 2025)) are retained. Detected categories undergo semantic clustering and merging, resulting in 566 semantically refined categories. These categories are translated into Chinese, and the final dataset consists of 24,221 images and 66,258 high-quality region-text pairs. We provide examples of the proposed datasets in Appendix I and J. Together, these datasets provide a rigorous and comprehensive assessment for bilingual vision-language alignment models, supporting deeper evaluation of fine-grained understanding capability.